OpenStudy (anonymous):

find the x-intercepts of the parabola with vertex (4,-1) and y-intercept (0,15) write your answer in this form:(x1,y1),(x2,y2) f necessary, round to the nearest Hundredth

OpenStudy (campbell_st):

well the vertex form of a parabola is \[y=a(x - h)^2 + k\] where (h, k) is the vertex. so substitute you vertex values, h = 4 and k = -1 into the equation. tThe y-intercept is on the parabola... so to find the value of a you need to substitute x = 0 and y = 15 and solve for a. then you can rewrite the equation of the parabola in standard form that's the 1st part hope it makes sense

OpenStudy (campbell_st):

once you have the equation in standard form you may need to use the general quadratic formula for the x intercepts

OpenStudy (anonymous):

soo would it be y=x^2-8x+15

OpenStudy (campbell_st):

that's the correct equation... and the equation can be factored.

OpenStudy (anonymous):

its wrong

OpenStudy (campbell_st):

so what did you get for the x-intercepts

OpenStudy (anonymous):

-4

OpenStudy (campbell_st):

the equation is \[y=x^2 - 8x + 15~~~~then~~~~y=(x -5)(x-3)\]

OpenStudy (campbell_st):

so the intercepts are

OpenStudy (campbell_st):

where y = 0 so solve x - 5 = 0 and x - 3 = 0

OpenStudy (anonymous):

am confused

OpenStudy (campbell_st):

ok.... so the x-intercepts these are points on the x-axis so y = 0 if you have a parabola to find the x-intercepts let y = 0 and solve for x so \[0 = x^2 - 8x + 15~~~or~~~0 =(x -5)(x -3)\] if either of the factors (x -5) or (x -3) are equal to zero the equation is equal to zero... so then you need to let the factors equal zero and solve for x so solve x - 5 = 0 and x - 3 = 0 what would you have...?

OpenStudy (campbell_st):

the other thing to think about is that the x-intercepts are the same distance from the x value in the vertex .

OpenStudy (anonymous):

am go fail this test

OpenStudy (campbell_st):

so if you the solve the 2 equations what do you get x - 5 = 0 and x - 3 = 0

OpenStudy (campbell_st):

these values are the x-intercepts

OpenStudy (anonymous):

x=5,3

OpenStudy (campbell_st):

correct so the intercepts are at (3,0) and (5, 0) the vertex has an x value of x = 4... the intercepts are the same distance from this value

OpenStudy (anonymous):

thank u soo much

OpenStudy (campbell_st):

yw

OpenStudy (anonymous):

it was wrong

OpenStudy (campbell_st):

so did you write the intercepts as (3, 0) and (5, 0)..?

OpenStudy (anonymous):

yea

OpenStudy (campbell_st):

here is a graph of the parabola you gave... and it meets all the conditions... can you check the vertex is (4, -1) and y-intercept is y = 15 or (0. 15)